# Security and Privacy on Generative Data in AIGC: A Survey

### v2

@@
Abstract
The advent of artificial intelligence-generated content (AIGC) represents a pivotal moment in the evolution of information technology. With AIGC, it can be effortless to generate high-quality data that is challenging for the public to distinguish. Nevertheless, the proliferation of generative data across cyberspace brings security and privacy issues, including privacy leakages of individuals and media forgery for fraudulent purposes. Consequently, both academia and industry begin to emphasize the trustworthiness of generative data, successively providing a series of countermeasures for security and privacy. In this survey, we systematically review the security and privacy on generative data in AIGC, particularly for the first time analyzing them from the perspective of information security properties. Specifically, we reveal the successful experiences of state-of-the-art countermeasures in terms of the foundational properties of privacy, controllability, authenticity, and compliance, respectively. Finally, we summarize the open challenges and potential exploration directions from each of theses properties.
Index Terms:
@@ Abstract
I Introduction
I-A Background
Artificial intelligence-generated content (AIGC) emerges as a novel generation paradigm for the production, manipulation, and modification of data. It utilizes advanced artificial intelligence (AI) technologies to automatically generate high-quality data at a rapid pace, including images, videos, text, audio, and graphics. With the powerful generative ability, AIGC can save time and unleash creativity, which are often challenging to achieve with professionally generated content (PGC) and user-generated content (UGC). Such progress in data creation can drive the emergence of innovative industries, particularly Metaverse [1], where digital and physical worlds converge. Early AIGC is limited by the algorithmic efficiency, hardware performance, and data scale, hindering the ability to fulfill optimal creation tasks. With the iterative updates of generative structures, notably generative adversarial networks (GANs) [2], AIGC has witnessed significant breakthroughs, generating realistic data that is often indistinguishable by humans from real data. In the generation of visual content, NVIDIA released StyleGAN [3] in 2018, which enables the controllable generation of high-resolution images and has undergone several upgrades. The subsequent year, DeepMind released DVD-GAN [4], which is designed for continuous video generation and exhibits great efficacy in complex data domains. Recently, diffusion models (DMs) [5] show more refined and novel image generation via the incremental noise addition. Guided by language models, DMs can improve the semantic coherence between input prompts and generated images. Excellent diffusion-based products, e.g., Stable Diffusion [^1] https://stability.ai/stablediffusion, Midjourney [^2] https://www.midjourney.com/, and Make-A-Video [^3] https://makeavideo.studio/, are capable of generating visually realistic images or videos that meet the requirements of diverse textual prompts. In the generation of language content, more attention is focused on ChatGPT, which reached 1.76 billion visits in May 2023. Trained on a large-scale text dataset, ChatGPT exhibits impressive performance in various contexts, including human-computer interaction and dialogues. For instance, researchers released LeanDojo [6], an open-source mathematical proof platform based on ChatGPT, providing toolkits, benchmarks, and models to tackle complex proof of formulas in an interactive environment. The integration of ChatGPT into the Bing search engine enhances search experiences, enabling users to effortlessly access comprehensive information. This powerful multi-purpose adaptability further exemplify the possibilities for humanity to achieve artificial general intelligence (AGI). Overall, compared to the PGC and UGC, AIGC demonstrates more advantages in data creation. AIGC possesses the ability to swiftly produce high-quality content while catering to personalized demands from users. As AI technology continues to advancements, the generative capability of AIGC is growing rapidly, promoting increased social productivity and economic value.
I-B Motivation
A large amount of generative data floods cyberspace, further enriching the diversity and abundance of online content. These generative data encompass multimodal information, which can be observed in various domains, e.g., news reporting, computer games, and social sharing. According to the Gartner’s report, AIGC will be anticipated to account for over 10% of all data creation in 2025 [7]. However, the proliferation of generative data also poses security and privacy issues. Firstly, generative data can expose individual privacy content by replicating training data. Generative models rely on large-scale data, which includes private content, e.g., faces, addresses, and emails. Existing works have demonstrated the memorization capabilities of large generative models [8, 9], leading to the potential replication of all or parts of the training data. This means that the generative data may also contain sensitive content which present in the training data. With specific prompts, GPT-2 can output personal information, including the name, the address, and the e-mail address [8]. An alarming study [9] revealed that Google’s Imagen can be prompted to output real-person photos, posing a significant threat to individual privacy. Therefore, it is necessary to hinder the generation of data containing privacy content. Secondly, generative data used for malicious purposes often involves false information, which can deceive the public, posing potential threats to both society and individuals. Recently, a false tweet about an explosion near the Pentagon went viral on social media, fooling many authoritative media sources and triggering fluctuations in the US stock market [10]. Moreover, the mature DeepFake technologies allow for the creation of convincing fake personal videos, which can be used to maliciously fabricate celebrity events [11]. The difficulty in discerning authenticity exposes the public to believing such content, resulting in severe damage to the reputations of celebrities. Thus, it is important to provide effective technology to confirm the authenticity of generative data. Meanwhile, generative data is required to have the controllability so that such potential threats can be proactively prevented. Thirdly, regulators around the world have further requirements for the compliance of generative data due to the critical implications of AIGC. Data protection regulatory authorities in Italy, Germany, France, Spain, and other countries have expressed concerns and initiated investigations into AIGC. In particular, China has taken a significant step by introducing the interim regulation on the management of generative artificial intelligence (AI) services [12]. This regulation encourages innovation in AIGC while mandating that generative data is non-toxic and factual. To adhere to the relevant regulations, it becomes crucial to ensure the compliance of generative data.
TABLE I: Comparisons of our work with existing surveys.
Reference Privacy Controllability Authenticity Compliance
Privacy in AIGC AIGC for Privacy Access Control Traceability Generative Detection Generative Attribution Non-toxicity Factuality
[13]
[14]
[15]
@@ I-B Motivation
I-C Comparisons with Existing Surveys
@@
Several works [13, 14, 15] have investigated the security and privacy in AIGC from different perspectives. Wang et al. [13] presents an in-depth survey of AIGC working principles, and roundly explored the taxonomy of security and privacy threats to AIGC. Meanwhile, they extensively reviewed solutions for intellectual property (IP) protection for AIGC models and generative data, with a focus on watermarking. Yet, they fail to provide countermeasures for other threats such as the utilization of non-compliant data.
@@
Chen et al. [14] discussed three main concerns for promoting responsible AIGC, including 1) privacy, 2) bias, toxicity, misinformation, and 3) intellectual property. They summarized the issues and listed solutions related to existing AIGC products, e.g., ChatGPT, Midjourney, and Imagen. Nevertheless, they overlooked the importance of considering the authenticity of generative data in responsible AIGC.
@@
Chen et al. [15] summarized the AIGC technology and analyzed the security and privacy challenges in AIGC. Moreover, they explored the potential countermeasures with advanced technologies, involving privacy computing, blockchain, and beyond. However, they did not pay attention to the detection and access control of generative data.
Figure 1: The process of AIGC. Real data collected is used to train generative models. Then generative models produce generative data. Finally, generative data are further analyzed. For generative data, there are corresponding protection requirements of security and privacy at different stages, which can be divided into privacy, controllability, authenticity, and compliance.
The differences between our work and previous works are: (1) Our work is targeted at generative data rather than AIGC. Previous works also explored issues about privacy in data collection and security of models. Yet, these issues are generic to AI security, which has been discussed in some works [16, 17]. (2) Previous works presented the corresponding techniques in terms of specific issues but neglected the fact that the same technique can address multiple issues. For example, watermarking can be used to protect copyright while detecting false data. On the contrary, we discuss security and privacy from the fundamental properties of information security, whose presence are ensured through specific techniques. (3) We supplement security issues that are not discussed in previous works, including access control and generative detection. In addition, we explore the use of generative data to power the privacy protection of real data. Table I shows the comparisons of our work with existing surveys.
@@
The rest of the paper is organized as follows. Section II reviews the basic AIGC process and categorizes the security and privacy on generative data in AIGC. In Sections III to VI, we discuss the issues and review the corresponding solutions from the perspectives of privacy, controllability, authenticity, and compliance, respectively. We present our outlook and suggest some future directions in Section VII. Finally, we summarize our work in Section VIII.
II Overview
II-A Process of AIGC
@@
II-A1 Real Data for Training
The data used for training impacts the features and patterns learned by AIGC models. Therefore, high-quality data forms the cornerstone of AIGC technology. Data collection typically involves various open-source repositories, including public databases, social media platforms, and online forums. These diverse sources provide AIGC training with a large-scale and diverse dataset. After collection, data filtering is applied to ensure the data quality, which involves removing irrelevant data and balancing the dataset for unbiased training. Additionally, data preprocessing, data augmentation, and data privacy protection steps can be undertaken based on different tasks to further enhance the quality and security of the training data.
II-A2 Generative Model in Training
The obtained data is used to train generative models, which are often performed by a centralized server with powerful computational capabilities. During training, generative models learn patterns and features in the data to generate results with a similar distribution to real data. Popular generative model architectures include generative adversarial networks (GANs), variational autoencoders (VAEs), flow-based models (Flows), and diffusion models (DMs), each with its strengths and weaknesses. The choice of models depends on specific requirements of tasks, available data, and computational resources. It is also important to note that training generative models requires substantial computational resources. On this basis, model fine-tuning is the process of adapting the pre-trained large model to a new task or domain without retraining. It only adjusts model parameters by training appropriate amount of additional data.
II-A3 Generative Data
After generative models are trained, they can be utilized to produce data. During this stage, users typically provide an input condition, e.g., a question or a piece of text. Then the model starts outputting data based on the input condition. In the generation of language content, AIGC exhibits the capability to outpace human authors in rapidly generating high-quality text, e.g., codes and articles. Additionally, it can engage in conversational interactions akin to humans, assisting users with various tasks and inquiries. The efficiency of AIGC in content creation and human-like interactions revolutionize how information is produced and communicated. In the generation of visual content, AIGC harnesses the powerful generative capabilities of models like DMs, enabling the generation of new images with realistic quality. Moreover, AIGC holds potential for video generation, as they can simultaneously process multiple video frames automatically.
II-A4 Analysis for Generative Data
@@
Generative data needs to undergo a quality assessment to check its accuracy, consistency, and integrality. If the generative data lacks in certain aspects, it requires model adjustments to improve the quality of generative data. Additionally, analyzing the risks associated with generative data can identify potential hazards. For instance, it is required to analyze whether there is discriminatory content, false information, or misleading content. By promptly detecting and addressing these issues, the negative impact of generative data can be minimized.
II-B Security and Privacy on Generative Data
Figure 2: The subclassification of security and privacy on generative data.
@@ II-B2 Controllability
  • Tracebility: Generative data needs to support the tracking of the generation process and subsequent dissemination for monitoring any behavior involving security.
II-B3 Authenticity
@@ II-B4 Compliance
III Privacy on Generative Data
@@
Figure 3: The example of privacy in AIGC is from [9], and the example of AIGC for privacy is from [18].
III-A Privacy for AIGC
III-A1 Threats to Privacy
@@
AIGC service providers proactively collect individual data on various platforms to construct giant datasets for enhancing the quality of generative data. However, the training data contains sensitive information about individuals, which is highly susceptible to privacy leakage. A study shows that the larger the amount of training data, the higher the privacy risk will result [19]. Specifically, in the training, private data can easily be memorized in model weights. In the interaction with users, generative data may replicate the training data, which poses a potential privacy threat. Such data replication is defined as an under-explored failure mode of overfitting, which exists in various generative models [20]. In language generative models, Carlini et al. [8] extracted training data by querying large language models (LLMs). The experiments employ GPT-2 as a demonstration to extract sensitive individual information, e.g., name, email, and phone number. Tirumala et al. [21] empirically studied the memorization dynamics over language model training, and demonstrated that larger models memorize faster. Nicholas et al. [22] described three log-linear relationships to quantify the extent to which LLMs memorize training data under different model scales, times of sample replications, and number of tokens. Their experiments indicated that memory in LLMs is more prevalent than previously believed and that memorization scales log-linearly with model size. In vision generative models, the general view is that generative adversarial networks (GANs) tend not to memorize training data under normal training settings [23]. However, Feng et al. [24] showed experimentally that GANs replication percentage decays exponentially with respect to dataset size and complexity. Stronger memory exists in diffusion models [25, 9, 26]. Carlini et al. [9] illustrated the ability of diffusion models to memorize individual images from the training data and can reproduce them at generation time. Compared with GANs, diffusion models are less private. Somepalli et al. [26] proposed image retrieval frameworks to demonstrate that the generated images by diffusion models are simple combinations of the foreground and background objects of the training dataset. As shown in Fig. 4, diffusion models just create semantically rather than pixel-wise objects identical to original images.
@@
Figure 4: Generated samples by Stable Diffusion just replicate training data via piecing together foreground and background objects in training data [26].
III-A2 Countermeasures to Privacy
Researchers have suggested some available solutions to mitigate data replication for privacy protection. Bai et al. [27] proposed memorization rejection in the training loss, which abandons generative data that are near-duplicates of training data. Deduplicating training datasets is also a possible option. OpenAI has verified its effectiveness using the distributed nearest neighbor search on Dall-E2 [28]. Nikhil et al. [29] studied a variety of LLMs and showed that the likelihood of a duplicated text sequence appearing is correlated with the number of occurrences of that sequence in the training data. In addition, they also verified that replicated text sequences are greatly reduced when duplicates are eliminated. Differential privacy [30] is a recommended solution, which introduces noises during training to ensure the generative data is differentially private with respect to the training data. RDP-GAN [31] adds differential noises on the value of the loss function of a discriminator during training, which achieves a differentially private GAN. DPDMs [32] enhances privacy via differentially private stochastic gradient descent, which also allows the generative data to retain a high level of availability to support multiple vision tasks. Compared to DPDMs, Ghalebikesabi et al. [33] enabled to accurately train larger models and to achieve a high utility on more challenging datasets such as CIFAR-10. Detecting replicated content is a remedial solution, which detects whether the generative data is in training data and then determines whether to use it. Stability AI provides a tool [34] to support the identification of the replicated images. Somepalli et al. [26] developed image similarity metrics that are diverse on self-supervised learning and based on an image retrieval framework to search for copying behavior. Machine unlearning [35] can help generative models forget the private training data, which avoids the effort of re-training the model. Kumari et. al [36] fine-tuned diffusion models by modifying the sensitive training data so that the models forget already memorized images. Forget-Me-Not [37] is adapted as a lightweight model patch for Stable Diffusion. It effectively removes the concept of containing a specific identity and avoids generating any face photo with the identity.
III-B AIGC for Privacy

### v3

@@
Abstract.
The advent of artificial intelligence-generated content (AIGC) represents a pivotal moment in the evolution of information technology. With AIGC, it can be effortless to generate high-quality data that is challenging for the public to distinguish. Nevertheless, the proliferation of generative data across cyberspace brings security and privacy issues, including privacy leakages of individuals and media forgery for fraudulent purposes. Consequently, both academia and industry begin to emphasize the trustworthiness of generative data, successively providing a series of countermeasures for security and privacy. In this survey, we systematically review the security and privacy on generative data in AIGC, particularly for the first time analyzing them from the perspective of information security properties. Specifically, we reveal the successful experiences of state-of-the-art countermeasures in terms of the foundational properties of privacy, controllability, authenticity, and compliance, respectively. Finally, we show some representative benchmarks, present a statistical analysis, and summarize the potential exploration directions from each of theses properties.
@@ Abstract.
This work is supported in part by the National Key R&D Program of China under Grant number 2022YFB3103100, in part by the National Natural Science Foundation of China under grant numbers 62122032, U23B2023.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8copyright: acmlicensedjournal: CSURjournalyear: 2024journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3703626ccs: Security and privacy Social aspects of security and privacyccs: Security and privacy Privacy protections
1. Introduction
$1.$ Background
Artificial intelligence-generated content (AIGC) emerges as a novel generation paradigm for the production, manipulation, and modification of data. It utilizes advanced artificial intelligence (AI) technologies to automatically generate high-quality data at a rapid pace, including images, videos, text, audio, and graphics. With the powerful generative ability, AIGC can save time and unleash creativity, which are often challenging to achieve with professionally generated content (PGC) and user-generated content (UGC). Such progress in data creation can drive the emergence of innovative industries, particularly Metaverse (Wang et al., 2022b), where digital and physical worlds converge.
Early AIGC is limited by the algorithmic efficiency, hardware performance, and data scale, hindering the ability to fulfill optimal creation tasks. With the iterative updates of generative structures, notably generative adversarial networks (GANs) (Goodfellow et al., 2020), AIGC has witnessed significant breakthroughs, generating realistic data that is often indistinguishable by humans from real data.
In the generation of visual content, NVIDIA released StyleGAN (Karras et al., 2019) in 2018, which enables the controllable generation of high-resolution images and has undergone several upgrades. The subsequent year, DeepMind released DVD-GAN (Clark et al., 2019), which is designed for continuous video generation and exhibits great efficacy in complex data domains. Recently, diffusion models (DMs) (Ho et al., 2020) show more refined and novel image generation via the incremental noise addition. Guided by language models, DMs can improve the semantic coherence between input prompts and generated images. Excellent diffusion-based products, e.g., Stable Diffusion [^1] https://stability.ai/stablediffusion, Midjourney [^2] https://www.midjourney.com/, and Make-A-Video [^3] https://makeavideo.studio/, are capable of generating visually realistic images or videos that meet the requirements of diverse textual prompts.
Table 1. Comparisons of our work with existing surveys.
Ref. Privacy Controllability Authenticity Compliance
Privacy AIGC for Access Traceability Generative Generative Non-toxicity Factuality
in AIGC Privacy Control Detection Attribution
(Wang et al., 2023c)
(Lyu et al., 2023a)
(Chen et al., 2023b)
@@ 1. Introduction
In the generation of language content, more attention is focused on ChatGPT, which reached 1.76 billion visits in May 2023. Trained on a large-scale text dataset, ChatGPT exhibits impressive performance in various contexts, including human-computer interaction and dialogues. For instance, researchers released LeanDojo (Yang et al., 2024a), an open-source mathematical proof platform based on ChatGPT, providing toolkits, benchmarks, and models to tackle complex proof of formulas in an interactive environment. The integration of ChatGPT into the Bing search engine enhances search experiences, enabling users to effortlessly access comprehensive information. This powerful multi-purpose adaptability further exemplifies the possibilities for humanity to achieve artificial general intelligence (AGI).
Overall, compared to the PGC and UGC, AIGC demonstrates more advantages in data creation. AIGC possesses the ability to swiftly produce high-quality content while catering to personalized demands from users. As AI technology continues to advancements, the generative capability of AIGC is growing rapidly, promoting increased social productivity and economic value.
1.2. Motivation
A large amount of generative data floods cyberspace, further enriching the diversity and abundance of online content. These generative data encompass multimodal information, which can be observed in various domains, e.g., news reporting, computer games, and social sharing. According to the Gartner’s report, AIGC will be anticipated to account for over 10% of all data creation in 2025. However, the proliferation of generative data also poses security and privacy issues.
Firstly, generative data can expose individual privacy content by replicating training data. Generative models rely on large-scale data, which includes private content, e.g., faces, addresses, and emails. Existing works have demonstrated the memorization capabilities of large generative models (Carlini et al., 2021, 2023a), leading to the potential replication of all or parts of the training data. This means that the generative data may also contain sensitive content which present in the training data. With specific prompts, GPT-2 can output personal information, including the name, the address, and the e-mail address (Carlini et al., 2021). An alarming study (Carlini et al., 2023a) revealed that Google’s Imagen can be prompted to output real-person photos, posing a significant threat to individual privacy. Therefore, it is necessary to hinder the generation of data containing privacy content.
Secondly, generative data used for malicious purposes often involves false information, which can deceive the public, posing potential threats to both society and individuals. Recently, a false tweet about an explosion near the Pentagon went viral on social media, fooling many authoritative media sources and triggering fluctuations in the US stock market (Pen, 2023). Moreover, the mature DeepFake technologies allow for the creation of convincing fake personal videos, which can be used to maliciously fabricate celebrity events (Korshunov and Marcel, 2018). The difficulty in discerning authenticity exposes the public to believing such content, resulting in severe damage to the reputations of celebrities. Thus, it is important to provide effective technology to confirm the authenticity of generative data. Meanwhile, generative data is required to have the controllability so that such potential threats can be proactively prevented.
Thirdly, regulators around the world have further requirements for the compliance of generative data due to the critical implications of AIGC. Data protection regulatory authorities in Italy, Germany, France, Spain, and other countries have expressed concerns and initiated investigations into AIGC. In particular, China has taken a significant step by introducing the interim regulation on the management of generative artificial intelligence (AI) services (int, 2023). This regulation encourages innovation in AIGC while mandating that generative data is non-toxic and factual. To adhere to the relevant regulations, it becomes crucial to ensure the compliance of generative data.
1.3. Comparisons with Existing Surveys
Several works (Wang et al., 2023c; Lyu et al., 2023a; Chen et al., 2023b) have investigated the security and privacy in AIGC from different perspectives.
@@ 1.3. Comparisons with Existing Surveys
Wang et al. (Wang et al., 2023c) presents an in-depth survey of AIGC working principles, and roundly explored the taxonomy of security and privacy threats to AIGC. Meanwhile, they extensively reviewed solutions for intellectual property (IP) protection for AIGC models and generative data, with a focus on watermarking. Yet, they fail to provide countermeasures for other threats such as the utilization of non-compliant data.
@@
Chen et al. (Lyu et al., 2023a) discussed three main concerns for promoting responsible AIGC, including 1) privacy, 2) bias, toxicity, misinformation, and 3) intellectual property. They summarized the issues and listed solutions related to existing AIGC products, e.g., ChatGPT, Midjourney, and Imagen. Nevertheless, they overlooked the importance of considering the authenticity of generative data in responsible AIGC.
@@
Chen et al. (Chen et al., 2023b) summarized the AIGC technology and analyzed the security and privacy challenges in AIGC. Moreover, they explored the potential countermeasures with advanced technologies, involving privacy computing, blockchain, and beyond. However, they did not pay attention to the detection and access control of generative data.
The differences between our work and previous works are:
  • Our work is targeted at generative data rather than AIGC. Previous works also explored issues about privacy in data collection and security of models. Yet, these issues are generic to AI security, which has been discussed in some works (Hu et al., 2021; Liu et al., 2021).
  • Previous works presented the corresponding techniques in terms of specific issues of privacy and security, whereas the issues cannot be enumerated in full. On the other hand, we discuss security and privacy from the fundamental properties of information security, which can cover the almost all of the issues.
  • We supplement security issues that are not discussed in previous works, including access control and generative detection. In addition, we explore the use of generative data to power the privacy protection of real data. Table 1 shows the comparisons of our work with existing surveys.
@@
The rest of the paper is organized as follows. Section 2 reviews the basic AIGC process and categorizes the security and privacy on generative data in AIGC. In Sections 3 to 6, we discuss the issues and review the corresponding solutions from the perspectives of privacy, controllability, authenticity, and compliance, respectively. We present some benchmarks and suggested future directions in Section 7 and Section 8. Finally, we summarize our work in Section 9.
Figure 1. The process of AIGC. Real data collected is used to train generative models. Then generative models produce generative data. Finally, generative data are further analyzed. For generative data, there are corresponding protection requirements of security and privacy at different stages, which can be divided into privacy, controllability, authenticity, and compliance.
2. Overview
2.1. Process of AIGC
@@
2.1.1. Real Data for Training
The data used for training impacts the features and patterns learned by AIGC models. Therefore, high-quality data forms the cornerstone of AIGC technology. Data collection typically involves various open-source repositories, including public databases, social media platforms, and online forums. These diverse sources provide AIGC training with a large-scale and diverse dataset.
After collection, data filtering is applied to ensure the data quality, which involves removing irrelevant data and balancing the dataset for unbiased training. Additionally, data preprocessing, data augmentation, and data privacy protection steps can be undertaken based on different tasks to further enhance the quality and security of the training data.
2.1.2. Generative Model in Training
The obtained data is used to train generative models, which are often performed by a centralized server with powerful computational capabilities. During training, generative models learn patterns and features in the data to generate results with a similar distribution to real data. Popular generative model architectures include generative adversarial networks (GANs), variational autoencoders (VAEs), flow-based models (Flows), and diffusion models (DMs), each with its strengths and weaknesses. The choice of models depends on specific requirements of tasks, available data, and computational resources.
It is also important to note that training generative models requires substantial computational resources. On this basis, model fine-tuning is the process of adapting the pre-trained large model to a new task or domain without retraining. It only adjusts model parameters by training appropriate amount of additional data.
2.1.3. Generative Data
After generative models are trained, they can be utilized to produce data. During this stage, users typically provide an input condition, e.g., a question or a piece of text. Then the model starts outputting data based on the input condition.
In the generation of language content, AIGC exhibits the capability to outpace human authors in rapidly generating high-quality text, e.g., codes and articles. Additionally, it can engage in conversational interactions akin to humans, assisting users with various tasks and inquiries. The efficiency of AIGC in content creation and human-like interactions revolutionize how information is produced and communicated.
In the generation of visual content, AIGC harnesses the powerful generative capabilities of models like DMs, enabling the generation of new images with realistic quality. Moreover, AIGC holds potential for video generation, as they can simultaneously process multiple video frames automatically.
2.1.4. Analysis for Generative Data
@@
Generative data needs to undergo a quality assessment to check its accuracy, consistency, and integrality. If the generative data lacks in certain aspects, it requires model adjustments to improve the quality of generative data.
Additionally, analyzing the risks associated with generative data can identify potential hazards. For instance, it is required to analyze whether there is discriminatory content, false information, or misleading content. By promptly detecting and addressing these issues, the negative impact of generative data can be minimized.
$2.$ Security and Privacy on Generative Data
Figure 2. The subclassification of security and privacy on generative data.
@@ 2.2.2. Controllability
  • Traceability: Generative data needs to support the tracking of the generation process and subsequent dissemination to monitor any behavior involving security for accountability.
2.2.3. Authenticity
@@ 2.2.4. Compliance
3. Privacy on Generative Data
@@
Figure 3. The example of privacy in AIGC is from (Carlini et al., 2023a), and the example of AIGC for privacy is from (Chen et al., 2021).
3.1. Privacy for AIGC
3.1.1. Threats to Privacy
AIGC service providers proactively collect individual data on various platforms to construct giant datasets for enhancing the quality of generative data. However, the training data contains sensitive information about individuals, which is highly susceptible to privacy leakage. A study shows that the larger the amount of training data, the higher the privacy risk will result (Plant et al., 2022). Specifically, in the training, private data can easily be memorized in model weights. In the interaction with users, generative data may replicate the training data, which poses a potential privacy threat. Such data replication is defined as an under-explored failure mode of overfitting, which exists in various generative models (Meehan et al., 2020).
In language generative models, Carlini et al. (Carlini et al., 2021) extracted training data by querying large language models (LLMs). The experiments employ GPT-2 as a demonstration to extract sensitive individual information, e.g., name, email, and phone number. Tirumala et al. (Tirumala et al., 2022) empirically studied the memorization dynamics over language model training, and demonstrated that larger models memorize faster. Nicholas et al. (Carlini et al., 2023b) described three log-linear relationships to quantify the extent to which LLMs memorize training data under different model scales, times of sample replications, and number of tokens. Their experiments indicated that memory in LLMs is more prevalent than previously believed and that memorization scales log-linearly with model size.
@@ 3.1.1. Threats to Privacy
In vision generative models, the general view is that generative adversarial networks (GANs) tend not to memorize training data under normal training settings (Webster et al., 2019). However, Feng et al. (Feng et al., 2021) showed experimentally that GANs replication percentage decays exponentially with respect to dataset size and complexity. Stronger memory exists in diffusion models (Webster, 2023; Carlini et al., 2023a; Somepalli et al., 2023). Carlini et al. (Carlini et al., 2023a) illustrated the ability of diffusion models to memorize individual images from the training data and can reproduce them at generation time. Unlike DMs that accept training data as direct input, generators for GANs are trained using only indirect information (i.e., gradients from the discriminator) about the training data. Therefore, GANs are more private. Somepalli et al. (Somepalli et al., 2023) proposed image retrieval frameworks to demonstrate that the generated images by diffusion models are simple combinations of the foreground and background objects of the training dataset. As shown in Fig. 4, diffusion models just create semantically rather than pixel-wise objects identical to original images.
@@
Figure 4. Generated samples by Stable Diffusion just replicate training data via piecing together foreground and background objects in training data (Somepalli et al., 2023).
3.1.2. Countermeasures to Privacy
Researchers have suggested some available solutions to mitigate data replication for privacy protection. Bai et al. (Bai et al., 2022) proposed memorization rejection in the training loss, which abandons generative data that are near-duplicates of training data.
Deduplicating training datasets is also a possible option. OpenAI has verified its effectiveness using the distributed nearest neighbor search on Dall-E2. Nikhil et al. (Kandpal et al., 2022) studied a variety of LLMs and showed that the likelihood of a duplicated text sequence appearing is correlated with the number of occurrences of that sequence in the training data. In addition, they also verified that replicated text sequences are greatly reduced when duplicates are eliminated.
Differential privacy (Abadi et al., 2016) is a recommended solution, which introduces noises during training to ensure the generative data is differentially private with respect to the training data. RDP-GAN (Ma et al., 2023b) adds differential noises on the value of the loss function of a discriminator during training, which achieves a differentially private GAN. DPDMs (Dockhorn et al., 2022) enhances privacy via differentially private stochastic gradient descent, which also allows the generative data to retain a high level of availability to support multiple vision tasks. Compared to DPDMs, Ghalebikesabi et al. (Ghalebikesabi et al., 2023) enabled to accurately train larger models and to achieve a high utility on more challenging datasets such as CIFAR-10. PrivImage (Li et al., 2024a) was proposed based on that the distribution of public data used should be similar to the distribution of sensitive data. Therefore, PrivImage elaborately selects similar distribution pre-training data, facilitating the efficient creation of DP datasets with high fidelity and utility. Unlike existing work that only considers DM as a regular deep model, dp-promise (Wang et al., 2024a) was the first work to implement (approximate) DP using DM noise, It employs a two-stage DM training process to reduce the overall noise injection, effectively achieving the privacy-utility tradeoff.
Detecting replicated content is a remedial solution, which detects whether the generative data is in training data and then determines whether to use it. Stability AI provides a tool (Bea, 2022) to support the identification of the replicated images. Somepalli et al. (Somepalli et al., 2023) developed image similarity metrics that are diverse on self-supervised learning and based on an image retrieval framework to search for copying behavior.
Machine unlearning (Bourtoule et al., 2021) can help generative models forget the private training data, which avoids the effort of re-training the model. Kumari et. al (Kumari et al., 2023) fine-tuned diffusion models by modifying the sensitive training data so that the models forget already memorized images. Forget-Me-Not (Zhang et al., 2024c) is adapted as a lightweight model patch for Stable Diffusion. It effectively removes the concept of containing a specific identity and avoids generating any face photo with the identity.
3.2. AIGC for Privacy